OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. B = 86°, b = 21, c = 18

OpenStudy (anonymous):

\[21\div \sin c =18\div \sin \left( 20 \right)\] first step

OpenStudy (anonymous):

then cross multiply the 2 proportions giving you.... \[21\sin c=18\sin 20\]

OpenStudy (anonymous):

then divide both sides by 20 and you get 0.293 since the value is less than one the answer is 2 triangles